OpenStudy (anonymous):

solve using elimination- 7. 3x+3y=27 x-3y=-11 9.2x+4y=22 2x-2y=-8 11.5x-4=0 3x+y=24

OpenStudy (campbell_st):

for question 7, if you add the 2 equations you eliminate y and have an equation in x to solve 3x + 3y = 27 + x - 3y = -11 --------------- 4x = 16 now solve for x. when you get the answer substitute the value into either equation to find y. hope it helps

OpenStudy (anonymous):

so x=4? @campbell_st

OpenStudy (campbell_st):

thats's correct.... now use either equation and x = 4 to find y

OpenStudy (anonymous):

A bicycle store costs $2400 per month to operate. The store pays an average of $60 per bike. The average selling price of each bike is $120. How many bikes must the store sell each month to break even?

OpenStudy (anonymous):

@campbell_st